OpenStudy (anonymous):

A body goes from 2 m to 12 m mark and back to 10 m mark in 3 seconds its average speed is ?

OpenStudy (anonymous):

 Notice the total distance travelled is 12m in 3 seconds

OpenStudy (anonymous):

You have to know that average speed in an interval is= total distance travelled/ total time

OpenStudy (anonymous):

which equals 12/3 =4m/s

OpenStudy (anonymous):

thnxs!!! i need ur more help

OpenStudy (anonymous):

A PARTICLE OF MASS 0.5 gm MOVING ALON ALONG X-AXIS IS LOCATED AT x1=15m atT1=5secs and at x2=33m at 13sec.Its velocity is:

OpenStudy (anonymous):

While finding velocity there are two things that you must find - the magnitude and the direction. The magnitude is simply given by Displacement/ time Initially mass was at 15 and after some time at 33. It is easy to note that mass was displaced by 33-15 m =18m Now it took the particle (13-5) = 8 seconds to do that So magnitude of velocity will be 18/8 m/s Now note that mass is moving along positive direction of x axis (initially it was at 15 later it is at 33 ) .So direction of velocity will be POSITIVE X AXIS .